OpenStudy (anonymous):

Given the following functions f(x) and g(x), solve f[g(7)] and select the correct answer below: f(x) = 4x + 21 g(x) = 2x + 2 16 49 85 784

OpenStudy (cwrw238):

g(7) = 2(7) + 2 = 16

OpenStudy (cwrw238):

oops - my bad thinking now you need to plug x = 16 into f(x)

OpenStudy (cwrw238):

4(16) + 21 = ?

OpenStudy (cwrw238):

- this gives you the value of f[g(7)]

OpenStudy (anonymous):

85

OpenStudy (anonymous):

so C

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

can you help with another ?

OpenStudy (cwrw238):

one more

OpenStudy (anonymous):

Given the functions f(x) = 7x + 13 and g(x) = x + 2, which of the following functions represents f[g(x)] correctly? f[g(x)] = 7x + 15 f[g(x)] = 7x + 27 f[g(x)] = 7x2 + 15 f[g(x)] = 7x2 + 27

OpenStudy (cwrw238):

you need to replace the x in f(x) - that is 7x + 13 - by g(x) which is x + 2 then simplify so we get 7(x + 2) + 13 - can you simplify this further?

OpenStudy (anonymous):

yes but how ?

OpenStudy (cwrw238):

7(x + 2) + 13 = 7*x + 7*2 + 13 = ?

OpenStudy (cwrw238):

you distribute the 7 over the x + 2

OpenStudy (cwrw238):

- that was the first step

OpenStudy (cwrw238):

so now its just simplifying 7x + 7*2 + 13

OpenStudy (anonymous):

is it A ?

OpenStudy (cwrw238):

no what is 7*2 ?

OpenStudy (cwrw238):

7 times 2

OpenStudy (cwrw238):

surely you know what 7 times 2 is?

OpenStudy (cwrw238):

right so we have 7x + 14 + 13 - now is just one add and you have the answer

OpenStudy (cwrw238):

add the 2 numbers

OpenStudy (cwrw238):

correct 7x + 27

OpenStudy (anonymous):

thanks

OpenStudy (cwrw238):

yw